OpenStudy (anonymous):

Name the property illustrated in this statement. 5 + (m + p) = 5 + (p + m) Additive Identity Property Additive Inverse Property Commutative Property Associative Property

Directrix (directrix):

5 + (m + p) = 5 + (p + m) @Ashja What do you think it is?

OpenStudy (anonymous):

@Directrix Commutative Property?

hero (hero):

Too bad we can't post the pictures directly

hero (hero):

Yes, correct @Ashja

hero (hero):

Commutative Property of Addition to be most specific

Directrix (directrix):

@Ashja Yes. A lot of people see the parenthesis and think "Distributive Property" but looking closely, 5 + (m + p) = 5 + (p + m), only the p and m exchanged places, and they are still grouped together.

OpenStudy (anonymous):

One more thing, is |-67 - 16|= 83?

OpenStudy (anonymous):

Or would it be -83?

hero (hero):

|| = absolute value = distance Distance is always positive

OpenStudy (anonymous):

Thanks!